Kyrie eleison, Lutheran (F–Mixolydian)

M: from Straßburg, 1524 after Christ • E: Norbert Südland, 2010 after Christ • Choral composition: April, 17^{th}, 2010 after Christ: Norbert Südland

[figure: 4 2; 4 2; Ch:; Ky-; ri-; e; e-; lei-; son.; All:; Pi-; ty; us,; oh; Lord.]

[figure: Ch:; Chris-; te; e-; lei-; son.; All:; Show; mer-; cy; to; us,; Christ.]

Kyrie eleison, Lutheran (G–Mixolydian)

M: from Straßburg, 1524 after Christ • E: Norbert Südland, 2010 after Christ • Choral composition: April, 17^{th}, 2010 after Christ: Norbert Südland

[figure: 4 2; 4 2; Ch:; Ky-; ri-; e; e-; lei-; son.; All:; Pi-; ty; us,; oh; Lord.]

[figure: Ch:; Chris-; te; e-; lei-; son.; All:; Show; mer-; cy; to; us,; Christ.]